OpenStudy (anonymous):

A triangle has an area of 72 square inches. If the base of the triangle has a length of 18 inches, what is the height of the triangle? Use the formula for the area of a triangle: Area = (base)(height) I will give a medal!!!!!!!!!!!

OpenStudy (eric_d):



OpenStudy (eric_d):

simplify it

OpenStudy (ahsome):

The formula of the area of a triangle is: \[A=\frac{bh}{2}\] Substitute the values \[72=\frac{18*h}{2}\] Solve from there. \[144=18*h\] \[h=144/18\] \[h=8\]

OpenStudy (eric_d):

@Dustinself42

OpenStudy (anonymous):

Thank both of you

OpenStudy (eric_d):

You're most welcome @Dustinself42

OpenStudy (ahsome):

@Dustinself42 No problem :)